OpenStudy (anonymous):

help!!! use the pythagorean throrem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 12,16,20

OpenStudy (experimentx):

get a calculator and see if 12^2 + 16^2 = 20^2

OpenStudy (anonymous):

so would the answer be yes or no

OpenStudy (experimentx):

Find out yourself http://www.wolframalpha.com/input/?i=12%5E2+%2B+16%5E2+%3D+20%5E2

OpenStudy (anonymous):

thanks budd!:)

OpenStudy (anonymous):

or recognize 12 - 16 - 20 as \(4\times 3\) , \(4\times 4\), \(4\times 5\) so you have the well known 3 - 4 - 5 right triangle in disguise